find real values for x and y so that each of the following equations is true:
(2x-1)-(3y+2)i=7+2i

Answer by robertb(5830)

The corresponding real parts have to be equal: 2x-1 = 7, or 2x = 8, or x = 4.
The corresponding imaginary parts have to be equal: -3y-2 = 2, or -3y = 4, or y = -4/3.